Cody ordered game tickets that cost $7.50 for children and $12.00 for adults. She ordered 8 more children's tickets than adults tickets. Her total bill was $138. How many of each type of ticker did she order?
: Cody ordered game tickets that cost $7.50 for children and $12.00 for adults. She ordered 8 more children's tickets than adults tickets. Her total bill was $138. How many of each type of ticker did she order?

Answer by galileo(7)

Let x = number of tickets ordered for children at $7.50 each
x-8 = number of tickets ordered for adults at $12.00 each
Equation :
7.5(x) + 12(x-8) = 138
7.5x + 12x - 96 = 138
19.5x - 96 = 138

19.5x = 138 + 96
19.5x = 234
x = 234/19.5
x = 12 ( tickets ordered for children )
x-8 = 12-8 = 4 ( tickets ordered for adults )